[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Nov 21, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Nov 22 03:13:56 UTC 2016


HSU Shift Log for Mon Nov 21 2016 22:13:50 GMT-0500 (EST)

Shift: R

Forecaster: D Brown

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: NRL TC web became inaccessible after about 02z.  
Around the same time, the FNMOC TC page that had been down for several days  
now has updated imagery.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161122/ca09570a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list